A buffer is prepared by adding 0.7387 g of NH4Cl (MM = 53.4915 g/mol) to 1.00 L of 0.2098 M NH3. Kb (NH3) = 1.8 x 10—5 . Calculate the pH.
1
Expert's answer
2020-03-24T14:01:20-0400

Solution:

m(NH4Cl) = 0.7387 g;

n(NH4Cl) = m(NH4Cl) / M(NH4Cl) = (0.7387 g) / (53.4915 g/mol) = 0.0138 mol;

c(NH4Cl) = n(NH4Cl) / V(NH4Cl) = (0.0138 mol) / (1.00 L) = 0.0138 mol / L = 0.0138 M.

c(NH4+) = 0.0138 M

c(NH3) = 0.2098 M


NH3 is a weak base: NH3 + H2O = NH4+ + OH-;

NH4Cl is a salt: NH4Cl → NH4+ + Cl-;

thus NH4+ is a "common ion":

NH3 + H2O = NH4+ + OH- 

Let's use an ICE (Initial, Change, Equilibrium) table:




The equlibrium constant (Kb) is expressed as:

Kb = [OH-] * [NH4+] / [NH3];

Kb = [x] * [0.0138 + x] / [0.2098 - x].

Approximation: ignore -x, +x terms: 

Kb = [x] * [0.0138] / [0.2098] = 0.0138x / 0.2098 = 0.06578x.

0.06578x = Kb = 1.8*10-5;

x = (1.8*10-5) / 0.06578 = 0.00027365 = 2.7365*10-4;

x= [OH-] = 2.7365*10-4.


pH = 14 - pOH;

pH = 14 - (-log[OH-]) = 14 - (log[2.7365*10-4]) = 14 - (3.56) = 10.44

pH = 10.44.


Answer: pH = 10.44